Karen left her house heading east and her sister left two hours later heading west at a rate 5 mph faster than Karen. 6 hours after Karen left she and her sister were 600 miles apart. How fast was Karen moving?
Answer by stanbon(75887)

Karen left her house heading east and
her sister left two hours later heading west at a rate 5 mph faster than Karen.
6 hours after Karen left she and her sister were 600 miles apart.
How fast was Karen moving?
---------------------
Karen DATA;
Time = 6 hrs ; rate = r mph ; Distance = 6r miles
------------------
Sister DATA:
Time = 4 hrs ; rate = r+5 mph ; Distance = 4(r+5) = 4r+20 miles
-------------------
EQUATION:
distance + distance = 600 miles
6r + 4r+20 = 600
10r = 580
r = 58 mph (Karen's rate)
r+5 = 63 mph (sister's rate)
==============
Cheers,
Stan H.